### v2

@@ G.1 Main proof
where the last inequality comes from the definition of $P^{*}(W_{k})$, as well as the definition of $P^{*}(W)$. Recall that convergence metric $E^{\texttt{TT}}_{K}$ defined in (87) is in the order of
$E^{\texttt{TT}}_{K}\geq$ $\ \Omega\left(\gamma^{3}\|W_{k}-W^{*}\|^{2}_{H(W_{k})}+\gamma^{2}\left\|P_{k}-P^{*}(W_{k})\right\|^{2}\right)$ (129)
$\geq$ $\ \Omega\left(\min\{H(W_{k})\}\gamma^{3}\|W_{k}-W^{*}\|^{2}+\gamma^{2}\left\|P_{k}-P^{*}(W_{k})\right\|^{2}\right)$
@@
$S^{\texttt{TT}}_{K}=\frac{1}{K}\sum_{k=0}^{K}\left\|\frac{G(P_{k})}{\sqrt{F(P_{k})}}\right\|^{2}_{\infty}\leq\frac{1}{K}\sum_{k=0}^{K}\left(2\|P_{k}-P^{*}(W_{k})\|^{2}+\frac{2}{\gamma^{2}}\left\|W_{k}-W^{*}\right\|^{2}\right)\leq O(E^{\texttt{TT}}_{K})$ (130)
where the last inequality holds if $\gamma$ is sufficiently large. Considering that, $E^{\texttt{TT}}_{K}-S^{\texttt{TT}}_{K}\geq\Omega(E^{\texttt{TT}}_{K})\geq 0$ and the conclusion is reached directly from Theorem 4.2. ∎
Appendix H Proof of Theorem H.1: Convergence of Analog GD
@@
$E_{K}^{\texttt{AGD}}:=\frac{1}{K}\sum_{k=0}^{K-1}\Big{(}\left\|\nabla f(W_{k})\odot F(W_{k})-|\nabla f(W_{k})|\odot G(W_{k})\right\|^{2}+\left\|\nabla f(W_{k})\right\|^{2}_{H(W_{k})}\Big{)}.$ (131)
@@
Theorem H.1 (Convergence of Analog GD).
Under Assumption 3.23.3, it holds that
@@
Further, if $H^{\texttt{ASGD}}_{\min}:=\min_{k\in[K]}\min\{Q_{+}(W_{k})Q_{-}(W_{k})\}>0$, it holds that
$\frac{1}{K}\sum_{k=0}^{K-1}\|\nabla f(W_{k})\|^{2}\leq\frac{2L(f(W_{0})-f^{*})F_{\max}^{2}}{KH^{\texttt{ASGD}}_{\min}}.$ (133)
Proof of Theorem H.1.
The $L$-smooth assumption (Assumption 3.2) implies that
@@
Define the saturation vector $H(W_{k})\in{\mathbb{R}}^{D}$ by
$H(W_{k}):=$ $\ {F(W_{k})^{\odot 2}-G(W_{k})^{\odot 2}}={(F(W_{k})+G(W_{k}))\odot(F(W_{k})-G(W_{k}))}$ (138)
@@
Noticing that $\|W_{k+1}-W_{k}\|^{2}=\alpha^{2}\|\nabla f(W_{k})\odot F(W_{k})-|\nabla f(W_{k})|\odot G(W_{k})\|^{2}$ and averaging for $k$ from 00 to $K-1$, we have
$E_{K}^{\texttt{AGD}}=$ $\ \frac{1}{K}\sum_{k=0}^{K-1}\Big{(}\left\|\nabla f(W_{k})\odot F(W_{k})-|\nabla f(W_{k})|\odot G(W_{k})\right\|^{2}+\left\|\nabla f(W_{k})\right\|^{2}_{H(W_{k})}\Big{)}$ (141)
@@
where the last inequality choose $\alpha=\frac{1}{2LF_{\max}}$.
Further, if the degree of saturation is bounded, (134)–(139) implies that
$\frac{\alpha H^{\texttt{AGD}}_{\min}}{2}\|\nabla f(W_{k})\|^{2}\leq\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}_{H(W_{k})}\leq f(W_{k})-f(W_{k+1}).$ (142)
Averaging (142) for $k$ from 00 to $K$ deduce that
$\frac{1}{K}\sum_{k=0}^{K-1}\|\nabla f(W_{k})\|^{2}\leq\frac{2(f(W_{0})-f(W_{K+1}))F_{\max}}{\alpha KH^{\texttt{AGD}}_{\min}}\leq\frac{2L(f(W_{0})-f^{*})F_{\max}^{2}}{KH^{\texttt{AGD}}_{\min}}$ (143)
@@
Appendix I Simulation Details and Additional Results
This section provides details about the experiments in Section 5. All simulation is performed under the PyTorch framework https://github.com/pytorch/pytorch. The analog training algorithms, including Analog SGD and Tiki-Taka, are provided by the open-source simulation toolkit AIHWKIT (Rasch et al., 2021), which has MIT license; see github.com/IBM/aihwkit.
Optimizer. The digital SGD optimizer is implemented by FloatingPointRPUConfig in AIHWKIT, which is equivalent to the SGD implemented in PyTorch. The Analog SGD is implemented by selecting SingleRPUConfig as configation, and Tiki-Taka optimizers are implemented by UnitCellRPUConfig with TransferCompound devices in AIHWKIT.
@@ Appendix I Simulation Details and Additional Results
I.1 Power and Exponential Response Functions
@@
It could be checked that the boundary of their dynamic ranges are $\tau^{\max}=\tau$ and $\tau^{\min}=-\tau$, while the symmetric point is 0, as required by Corollary 4.4. Figure 5 illustrates how the response functions change with different $\gamma_{\texttt{res}}$.
@@
I.2 Least squares problem
@@
The elements of $W^{*}$ are sampled from a Gaussian distribution with mean 00 and variance $\sigma^{2}_{W^{*}}$. Consider a matrix $A\in{\mathbb{R}}^{D_{\text{out}}\times D}$ of size $D=50$ and $D_{\text{out}}=100$ whose elements are sampled from a Gaussian distribution with variance $\sigma^{2}_{A}$. The label $b\in{\mathbb{R}}^{D_{\text{out}}}$ is generated by $b=AW^{*}$ where $W^{*}$ are sampled from a standard Gaussian distribution with $\sigma^{2}_{W^{*}}$. The response granularity $\Delta w_{\min}$=1e-4 while $\tau=3.5$. The maximum bit length is 8. The variance are set as $\sigma^{2}_{A}=1.00^{2}$, $\sigma^{2}_{W^{*}}=0.5^{2}$.
I.3 Classification problem
@@
The other setting follows the settings of AIHWKIT, including output noise (0.5 % of the quantization bin width), quantization and clipping (output range set 20, output noise 0.1, and input and output quantization to 8 bits). Noise and bound management techniques are used in Gokmen et al. (2017). A learnable scaling factor is set after each analog layer, which is updated using SGD.
3-FC / MNIST. Following the setting in Gokmen & Vlasov (2016), we train a model with 3 fully connected layers. The hidden sizes are 256 and 128. The activation functions are Sigmoid. The learning rates are $\alpha=0.1$ for Digital SGD, $\alpha=0.05,\beta=0.01$ for Analog SGD and Tiki-Taka. The batch size is 10 for all algorithms.
CNN / MNIST. We train a convolution neural network, which contains 2-convolutional layers, 2-max-pooling layers, and 2-fully connected layers. The activation functions are Tanh. The first two convolutional layers use 5$\times$5 kernels with 16 and 32 kernels, respectively. Each convolutional layer is followed by a subsampling layer implemented by the max pooling function over non-overlapping pooling windows of size 2 $\times$ 2. The output of the second pooling layer, consisting of 512 neuron activations, feeds into a fully connected layer consisting of 128 tanh neurons, which is then connected into a 10-way softmax output layer. The learning rates are set as $\alpha=0.1$ for Digital SGD, $\alpha=0.05,\beta=0.01$ for Analog SGD or Tiki-Taka. The batch size is 8 for all algorithms.
ResNet / CIFAR10 & CIFAR100. We train different models from the ResNet family, including ResNet18, 34, and 50. The base model is pre-trained on ImageNet dataset. The last fully connected layer is replaced by an analog layer. The learning rates are set as $\alpha=0.075$ for Digital SGD, $\alpha=0.075,\beta=0.01$ for Analog SGD or Tiki-Taka. Tiki-Taka adopts $\gamma=0.4$ unless stated otherwise. The batch size is 128 for all algorithms. Power response with $\gamma_{\texttt{res}}=3.0$ and $\tau=0.1$, and exponential response with $\gamma_{\texttt{res}}=3.0$ and $\tau=0.1$, are used in the simulations.
I.4 Additional performance on real datasets
@@
CIFAR10/CIFAR100 ResNet. We fine-tune three models from the ResNet family with different scales on CIFAR10/CIFAR100 datasets. The power response functions with the parameter $\gamma_{\text{res}}=3.0$ and $\tau=0.1$ is used, whose results are shown in Table 1 and 3, respectively. The results show that the Tiki-Taka outperforms Analog SGD by about $1.0\%$ in most of the cases in ResNet34/50, and the gap even reaches about $10.0\%$ for ResNet18 training on the CIFAR100 dataset. An interesting observation is that the gap is closer on the exponential response, at which time Analog SGD achieves a comparable and sometimes higher accuracy with Tiki-Taka. We speculate it happens because the asymmetric bias, coupled with the impact of a high-dimensional objective landscape, leads to a better local minimum, which is worthy of study in future work.
CIFAR10 CIFAR100
DSGD ASGD TT DSGD ASGD TT
ResNet18 95.43 $\pm$0.13 94.66 $\pm$0.11 94.70 $\pm$0.07 81.12 $\pm$0.25 73.55 $\pm$0.22 74.64 $\pm$0.24
ResNet34 96.48 $\pm$0.02 96.19 $\pm$0.04 96.24 $\pm$0.08 83.86 $\pm$0.12 78.10 $\pm$0.24 79.05 $\pm$0.21
ResNet50 96.57 $\pm$0.10 96.53 $\pm$0.06 96.40 $\pm$0.13 83.98 $\pm$0.11 81.40 $\pm$0.36 79.75 $\pm$0.10
@@
Table 3: Analog training with the exponential response for fine-tuning task on CIFAR10/100 datasets using models from ResNet family. The accuracy of the test set is reported. DSGD, ASGD, and TT represent Digital SGD, Analog SGD, Tiki-Taka, respectively.
@@
CIFAR10/CIFAR100 MobileNet. We fine-tune three models from the MobileNet family with different scales on CIFAR10/CIFAR100 datasets. The response function are set as the power response with the parameter $\gamma_{\text{res}}=4.0$ and $\tau=0.05$, whose results are shown in Table 4. In the simulations, the accuracy of Analog SGD drops significantly by about $10\%$ in most cases, while Tiki-Taka remains comparable to the Digital SGD with only a slight drop.
CIFAR10 CIFAR100
DSGD ASGD TT DSGD ASGD TT
MobileNetV2 95.28 $\pm$0.20 94.34 $\pm$0.27 95.05 $\pm$0.11 80.60 $\pm$0.18 63.41 $\pm$1.20 73.33 $\pm$0.94
MobileNetV3S 94.45 $\pm$0.10 80.66 $\pm$6.18 93.65 $\pm$0.24 78.94 $\pm$0.05 51.79 $\pm$1.05 71.14 $\pm$0.93
MobileNetV3L 95.95 $\pm$0.08 80.79 $\pm$2.97 95.39 $\pm$0.27 82.16 $\pm$0.26 66.80 $\pm$1.40 78.81 $\pm$0.52
@@
Table 4: Analog training with exponential response for fine-tuning task on CIFAR10/100 datasets using models from the MobileNet family. The accuracy of the test set is reported. DSGD, ASGD, and TT represent Digital SGD, Analog SGD, Tiki-Taka, respectively.
I.5 Ablation study on cycle variation
@@
To verify the conclusion of Theorem 2.1 that the error introduced by cycle variation is a higher-order term, we conduct a numerical simulation training on image classification task on MNIST dataset using FCN (FCN) or CNN (CNN) network. In the pulse update (2), the parameter $\sigma_{c}$ is varied from 10% to 120%, where the noise signal is already larger than the response function signal itself. The results are shown in Table 5. The results show that the test accuracy of both Analog SGD and Tiki-Taka is not significantly affected by the cycle variation, which complies with the theoretical analysis.
@@
DSGD ASGD TT DSGD ASGD TT
@@
Table 5: Test accuracy comparison under different cycle variation levels $\sigma_{c}$ on MNIST dataset. DSGD, ASGD, and TT represent Digital SGD, Analog SGD, Tiki-Taka, respectively
I.6 Ablation study on various response functions
We also train a FCN model on the MNIST dataset under various response functions. As shown in the figure, larger $\gamma_{\texttt{res}}$ leads to a steeper response function. The results are shown in Table 6. The accuracy $<15.00$ in the table implies that Analog SGD fails completely at all trials, which is close to random guess. The results show that Analog SGD works well only when the asymmetric is mild, i.e. $\gamma_{\texttt{res}}$ is small and $\tau$ is large, while Tiki-Taka outperforms Analog SGD and achieves comparable accuracy with Digital SGD.
@@ I.6 Ablation study on various response functions
ASGD TT ASGD TT
$\gamma_{\texttt{res}}=0.5$ $\tau=0.6$ 98.17 $\pm$0.05 96.01 $\pm$0.26 96.92 $\pm$0.19 $<$15.00 97.27 $\pm$0.07
$\tau=0.7$ 97.40 $\pm$0.15 97.05 $\pm$0.05 $<$15.00 97.39 $\pm$0.15
@@
$\gamma_{\texttt{res}}=1.0$ $\tau=0.6$ $<$15.00 97.39 $\pm$0.05 $<$15.00 97.46 $\pm$0.08
$\tau=0.7$ $<$15.00 97.33 $\pm$0.05 $<$15.00 97.49 $\pm$0.04
$\tau=0.8$ $<$15.00 97.34 $\pm$0.09 $<$15.00 97.25 $\pm$0.16
$\gamma_{\texttt{res}}=2.0$ $\tau=0.6$ $<$15.00 96.93 $\pm$0.15 $<$15.00 97.19 $\pm$0.16
$\tau=0.7$ $<$15.00 97.27 $\pm$0.02 $<$15.00 97.72 $\pm$0.07
$\tau=0.8$ $<$15.00 97.18 $\pm$0.04 $<$15.00 97.06 $\pm$0.10
Table 6: Test accuracy comparison under different response function parameters $\tau$ and $\gamma_{\texttt{res}}$ for FCN training on MNIST dataset with power or exponential response functions. DSGD, ASGD, and TT represent Digital SGD, Analog SGD, Tiki-Taka, respectively.
I.7 Ablation study on $\gamma$

### v3

@@ I.5 Proof of Corollary 1: Exact convergence of Tiki-Taka
where the last inequality comes from the definition of $P^{*}(W_{k})$, as well as the definition of $P^{*}(W)$. Recall that convergence metric $E^{\texttt{RL}}_{K}$ defined in (83) is in the order of
$E^{\texttt{RL}}_{K}\geq$ $\ \Omega\left(\gamma^{3}\|W_{k}-W^{*}\|^{2}_{R(W_{k})}+\gamma^{2}\left\|P_{k}-P^{*}(W_{k})\right\|^{2}\right)$ (124)
$\geq$ $\ \Omega\left(\min\{R(W_{k})\}\gamma^{3}\|W_{k}-W^{*}\|^{2}+\gamma^{2}\left\|P_{k}-P^{*}(W_{k})\right\|^{2}\right)$
@@
$S^{\texttt{RL}}_{K}=\frac{1}{K}\sum_{k=0}^{K}\left\|\frac{G(P_{k})}{\sqrt{F(P_{k})}}\right\|^{2}_{\infty}\leq\frac{1}{K}\sum_{k=0}^{K}\left(2\|P_{k}-P^{*}(W_{k})\|^{2}+\frac{2}{\gamma^{2}}\left\|W_{k}-W^{*}\right\|^{2}\right)\leq O(E^{\texttt{RL}}_{K})$ (125)
where the last inequality holds if $\gamma$ is sufficiently large. Considering that, $E^{\texttt{RL}}_{K}-S^{\texttt{RL}}_{K}\geq\Omega(E^{\texttt{RL}}_{K})\geq 0$ and the conclusion is reached directly from Theorem 3. ∎
Appendix J Proof of Theorem 6: Convergence of Analog GD
@@
[…]
@@
Theorem 6 (Convergence of Analog GD).
Under Assumption 12, it holds that
@@
Further, if $R^{\texttt{ASGD}}_{\min}:=\min_{k\in[K]}\min\{Q_{+}(W_{k})Q_{-}(W_{k})\}>0$, it holds that
$\frac{1}{K}\sum_{k=0}^{K-1}\|\nabla f(W_{k})\|^{2}\leq\frac{2L(f(W_{0})-f^{*})F_{\max}^{2}}{KR^{\texttt{ASGD}}_{\min}}.$ (128)
Proof of Theorem 6.
The $L$-smooth assumption (Assumption 1) implies that
@@
Define the saturation vector $R(W_{k})\in{\mathbb{R}}^{D}$ by
$R(W_{k}):=$ $\ {F(W_{k})^{\odot 2}-G(W_{k})^{\odot 2}}={(F(W_{k})+G(W_{k}))\odot(F(W_{k})-G(W_{k}))}$ (133)
@@
Noticing that $\|W_{k+1}-W_{k}\|^{2}=\alpha^{2}\|\nabla f(W_{k})\odot F(W_{k})-|\nabla f(W_{k})|\odot G(W_{k})\|^{2}$ and averaging for $k$ from 0 to K1K-1, we have
[…]
@@
where the last inequality choose $\alpha=\frac{1}{2LF_{\max}}$. Further, if the degree of saturation is bounded, (129)–(134) implies that
$\frac{\alpha R^{\texttt{AGD}}_{\min}}{2}\|\nabla f(W_{k})\|^{2}\leq\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}_{R(W_{k})}\leq f(W_{k})-f(W_{k+1}).$ (137)
Averaging (137) for $k$ from 0 to $K$ deduce that
$\frac{1}{K}\sum_{k=0}^{K-1}\|\nabla f(W_{k})\|^{2}\leq\frac{2(f(W_{0})-f(W_{K+1}))F_{\max}}{\alpha KR^{\texttt{AGD}}_{\min}}\leq\frac{2L(f(W_{0})-f^{*})F_{\max}^{2}}{KR^{\texttt{AGD}}_{\min}}$ (138)
@@
Appendix K Simulation Details and Additional Results
This section provides details about the experiments in Section 6. All simulation is performed under the PyTorch framework https://github.com/pytorch/pytorch. The analog training algorithms, including Analog SGD and Tiki-Taka, are provided by the open-source simulation toolkit AIHWKIT [44], which has MIT license; see github.com/IBM/aihwkit.
Optimizer. The digital SGD optimizer is implemented by FloatingPointRPUConfig in AIHWKIT, which is equivalent to the SGD implemented in PyTorch. The Analog SGD is implemented by selecting SingleRPUConfig as configuration and Tiki-Taka optimizers are implemented by UnitCellRPUConfig with TransferCompound devices in AIHWKIT.
@@
K.1 Power and Exponential Response Functions
@@
It could be checked that the boundary of their dynamic ranges are $\tau^{\max}=\tau$ and $\tau^{\min}=-\tau$, while the symmetric point is 0, as required by Corollary 1. Figure 5 illustrates how the response functions change with different $\gamma_{\texttt{res}}$.
@@
K.2 Least squares problem
@@
The elements of $W^{*}$ are sampled from a Gaussian distribution with mean 0 and variance $\sigma^{2}_{W^{*}}$. Consider a matrix $A\in{\mathbb{R}}^{D_{\text{out}}\times D}$ of size $D=50$ and $D_{\text{out}}=100$ whose elements are sampled from a Gaussian distribution with variance $\sigma^{2}_{A}$. The label $b\in{\mathbb{R}}^{D_{\text{out}}}$ is generated by $b=AW^{*}$ where $W^{*}$ are sampled from a standard Gaussian distribution with $\sigma^{2}_{W^{*}}$. The response granularity $\Delta w_{\min}$=1e-4 while $\tau=3.5$. The maximum bit length is 8. The variance are set as $\sigma^{2}_{A}=1.00^{2}$, $\sigma^{2}_{W^{*}}=0.5^{2}$.
K.3 Classification problem
@@
The other setting follows the settings of AIHWKIT, including output noise (0.5 % of the quantization bin width), quantization and clipping (output range set 20, output noise 0.1, and input and output quantization to 8 bits). Noise and bound management techniques are used in [71]. A learnable scaling factor is set after each analog layer, which is updated using SGD.
3-FC / MNIST. Following the setting in [16], we train a model with 3 fully connected layers. The hidden sizes are 256 and 128. The activation functions are Sigmoid. The learning rates are $\alpha=0.1$ for Digital SGD, $\alpha=0.05,\beta=0.01$ for Analog SGD and Tiki-Taka. The batch size is 10 for all algorithms. In Figure 4, the power response functions with $\gamma_{\text{res}}=0.5$ are used, and various $\tau$ are used as indicated in the legend.
CNN / MNIST. We train a convolution neural network, which contains 2-convolutional layers, 2-max-pooling layers, and 2-fully connected layers. The activation functions are Tanh. The first two convolutional layers use $\times5$ kernels with 16 and 32 kernels, respectively. Each convolutional layer is followed by a subsampling layer implemented by the max pooling function over non-overlapping pooling windows of size 2 $\times$ 2. The output of the second pooling layer, consisting of 512 neuron activations, feeds into a fully connected layer consisting of 128 tanh neurons, which is then connected into a 10-way softmax output layer. The learning rates are set as $\alpha=0.1$ for Digital SGD, $\alpha=0.05,\beta=0.01$ for Analog SGD are Residual Learning/Tiki-Taka. The batch size is 8 for all algorithms. In Figure 4, the power response functions with $\gamma_{\text{res}}=0.5$ are used, and various $\tau$ are used as indicated in the legend.
ResNet & MobileNet / CIFAR10 & CIFAR100. We train different models from the ResNet family, including ResNet18, 34, and 50. The base model is pre-trained on ImageNet dataset. The last fully connected layer is replaced by an analog layer. The learning rates are set as $\alpha=0.075$ for Digital SGD, $\alpha=0.075,\beta=0.01$ for Analog SGD, Residual Learning/Tiki-Taka, Tiki-Taka v2, and Residual Learningv2. Tiki-Taka adopts $\gamma=0.4$ unless stated otherwise. The batch size is 128 for all algorithms.
K.4 Additional performance on real datasets
@@
CIFAR10/CIFAR100 ResNet. We fine-tune three models from the ResNet family with different scales on CIFAR10/CIFAR100 datasets. The power response functions with $\gamma_{\text{res}}=3.0$ and $\tau=0.1$, and the exponential response functions with $\gamma_{\text{res}}=4.0$ and $\tau=0.1$ are used, whose results are shown in Table 1 and 3, respectively. The results show that the Tiki-Taka outperforms Analog SGD by about $1.0\%$ in most of the cases in ResNet34/50, and the gap even reaches about $10.0\%$ for ResNet18 training on the CIFAR100 dataset.
CIFAR10
DSGD ASGD TT/RL TTv2 RLv2
ResNet18 95.43 $\pm$0.13 84.47 $\pm$3.40 94.81 $\pm$0.09 95.31 $\pm$0.05 95.12 $\pm$0.14
ResNet34 96.48 $\pm$0.02 95.43 $\pm$0.12 96.29 $\pm$0.12 96.60 $\pm$0.05 96.42 $\pm$0.13
ResNet50 96.57 $\pm$0.10 94.36 $\pm$1.16 96.34 $\pm$0.04 96.63 $\pm$0.09 96.56 $\pm$0.08
CIFAR100
DSGD ASGD TT/RL TTv2 RLv2
ResNet18 81.12 $\pm$0.25 68.98 $\pm$1.01 76.17 $\pm$0.23 78.56 $\pm$0.29 79.83 $\pm$0.13
ResNet34 83.86 $\pm$0.12 78.98 $\pm$0.55 80.58 $\pm$0.11 81.81 $\pm$0.15 82.85 $\pm$0.19
ResNet50 83.98 $\pm$0.11 79.88 $\pm$1.26 80.80 $\pm$0.22 82.82 $\pm$0.33 83.90 $\pm$0.20
@@
Table 3: Fine-tuning ResNet models with the exponential response on CIFAR10/100 datasets. Test accuracy is reported. DSGD, ASGD, and $T$ represent Digital SGD, Analog SGD, Tiki-Taka, respectively.
@@
CIFAR10/CIFAR100 MobileNet. We fine-tune three MobileNet models with different scales on CIFAR10/CIFAR100 datasets. The response function is set as the power response with the parameter $\gamma_{\text{res}}=4.0$ and $\tau=0.05$, whose results are shown in Table 4. In the simulations, the accuracy of Analog SGD drops significantly by about $10\%$ in most cases, while Tiki-Taka remains comparable to the Digital SGD with only a slight drop.
CIFAR10
DSGD ASGD TT/RL TTv2 RLv2
MobileNetV2 95.28 $\pm$0.20 94.34 $\pm$0.27 95.05 $\pm$0.11 95.20 $\pm$0.14 95.26 $\pm$0.03
MobileNetV3S 94.45 $\pm$0.10 80.66 $\pm$6.18 93.65 $\pm$0.24 93.54 $\pm$0.06 93.79 $\pm$0.00
MobileNetV3L 95.95 $\pm$0.08 80.79 $\pm$2.97 95.39 $\pm$0.27 95.27 $\pm$0.09 95.33 $\pm$0.08
CIFAR100
DSGD ASGD TT/RL TTv2 RLv2
MobileNetV2 80.60 $\pm$0.18 63.41 $\pm$1.20 73.33 $\pm$0.94 78.41 $\pm$0.15 79.60 $\pm$0.10
MobileNetV3S 78.94 $\pm$0.05 51.79 $\pm$1.05 71.14 $\pm$0.93 74.51 $\pm$0.37 75.39 $\pm$0.00
MobileNetV3L 82.16 $\pm$0.26 66.80 $\pm$1.40 78.81 $\pm$0.52 79.56 $\pm$0.10 80.18 $\pm$0.07
@@
Table 4: Fine-tuning MobileNet models with power response on CIFAR10/100 datasets. Test accuracy is reported. DSGD, ASGD, and $T$ represent Digital SGD, Analog SGD, Tiki-Taka, respectively.
K.5 Ablation study on cycle variation
@@
To verify the conclusion of Theorem 4 that the error introduced by cycle variation is a higher-order term, we conduct a numerical simulation training on an image classification task on the MNIST dataset using FCN (FCN) or CNN (CNN) network. In the pulse update (25), the parameter $\sigma_{c}$ is varied from 10% to 120%, where the noise signal is already larger than the response function signal itself. The results are shown in Table 5. The results show that the test accuracy of both Analog SGD and Tiki-Taka is not significantly affected by the cycle variation, which complies with the theoretical analysis.
@@
DSGD ASGD $T$ DSGD ASGD $T$
@@
Table 5: Test accuracy comparison under different cycle variation levels $\sigma_{c}$ on MNIST dataset. DSGD, ASGD, and $T$ represent Digital SGD, Analog SGD, Tiki-Taka, respectively
K.6 Ablation study on various response functions
We also train a FCN model on the MNIST dataset under various response functions. As shown in the figure, larger $\gamma_{\texttt{res}}$ leads to a steeper response function. The results are shown in Table 6. The accuracy <15.00<15.00 in the table implies that Analog SGD fails completely at all trials, which is close to random guess. The results show that Analog SGD works well only when the asymmetric is mild, i.e. $\gamma_{\texttt{res}}$ is small and $\tau$ is large, while Tiki-Taka outperforms Analog SGD and achieves comparable accuracy with Digital SGD.
@@
ASGD TT/RL ASGD TT/RL
$\gamma_{\texttt{res}}=0.5$ $\tau=0.6$ 98.17 $\pm$0.05 96.01 $\pm$0.26 96.92 $\pm$0.19 <<15.00 97.27 $\pm$0.07
$\tau=0.7$ 97.40 $\pm$0.15 97.05 $\pm$0.05 <<15.00 97.39 $\pm$0.15
@@
$\gamma_{\texttt{res}}=1.0$ $\tau=0.6$ <<15.00 97.39 $\pm$0.05 <<15.00 97.46 $\pm$0.08
$\tau=0.7$ <<15.00 97.33 $\pm$0.05 <<15.00 97.49 $\pm$0.04
$\tau=0.8$ <<15.00 97.34 $\pm$0.09 <<15.00 97.25 $\pm$0.16
$\gamma_{\texttt{res}}=2.0$ $\tau=0.6$ <<15.00 96.93 $\pm$0.15 <<15.00 97.19 $\pm$0.16
$\tau=0.7$ <<15.00 97.27 $\pm$0.02 <<15.00 97.72 $\pm$0.07
$\tau=0.8$ <<15.00 97.18 $\pm$0.04 <<15.00 97.06 $\pm$0.10
Table 6: Test accuracy comparison under different response function parameters $\tau$ and $\gamma_{\texttt{res}}$ for FCN training on MNIST dataset with power or exponential response functions. DSGD, ASGD, and $T$ represent Digital SGD, Analog SGD, Tiki-Taka, respectively.
K.7 Ablation study on $\gamma$
@@
Appendix L Broader Impact
This paper focuses on developing a theoretical analysis for gradient-based training algorithms on a class of generic AIMC hardware, which can be leveraged to boost both energy and computational efficiency of training. While such efficiency gains could, in principle, enable broader and potentially unintended uses of machine learning models, we do not identify any specific societal risks that need to be highlighted in this context.